# Neural Honeytrace: A Robust Plug-and-Play Watermarking Framework against Model Extraction Attacks

### v3

@@ Appendix A Additional Experimental Results
A.1 Neural Honeytrace for Generative Models
Compared to classification models, the output space of generative models such as GAN and diffusion models has higher dimensions. According to our watermark transmission model, larger output spaces lead to better watermark transmission. Therefore, we evaluate Neural Honeytrace for Denoising Diffusion Probabilistic Models [53] (DDPM) trained on CIFAR-10. DDPM trains a U-Net model to predict the noise added in the current step and generates an image by iterative denoising:
@@ A.1 Neural Honeytrace for Generative Models
A.2 Overhead of Different Defenses
Fig. 11 compares the defense overhead of different triggerable watermarking methods for protecting target models trained on various datasets. During the training phase, previous triggerable watermarking methods introduce additional time costs due to modifications in the training process. For example, EWE[16] adds regularization terms to the loss function, increasing the time cost of each forward-backward pass. Composite Backdoor [18] and MEA-Defender [20] expand the training dataset with generated data. In contrast, Neural Honeytrace is training-free and can be directly applied to the target model without training.
A.3 Additional Ablation Study
In line with Tab. VI and Fig. 7, we provide ablation study results on CIFAR-100 in Tab. X and Fig. 12.
Different Query Datasets. In Tab.X, we compare the performance of Neural Honeytrace on a target model trained on CIFAR-100 when attackers use CIFAR-10, CIFAR-100, and TinyImageNet as surrogate datasets, respectively. The experimental results show that for out-of-distribution surrogate datasets (CIFAR-10 and TinyImageNet), attackers achieve similar extraction accuracy, and Neural Honeytrace maintains high watermark success rates. However, when using the same training dataset as the surrogate (CIFAR-100), attackers achieve higher extraction accuracy, while the watermark success rate of decreases.
The other columns in Fig. 12 show the effectiveness of the three hyperparameters, , of Neural Honeytrace. These hyperparameters are used to balance the model availability and the watermark success rate. According to Eq. 7, Eq. 9, and Eq. 10, intuitively, larger and smaller leads to stronger watermarks but lower protected accuracy, which can also be observed in Fig. 12. Therefore, given a test dataset and the acceptable maximum drop in accuracy, model owners can identify appropriate hyperparameters.
Test-time Backdoor Detection. Fig.13 presents additional results evaluating existing watermarking strategies against the test-time backdoor detection method BTI[34]. As depicted, BTI fails to effectively detect watermarks embedded in surrogate models for several reasons: (1) For EWE, the watermark success rate is significantly lower compared to conventional backdoor attacks, making it challenging to search for triggers. (2) For MEA-Defender, the composite trigger’s size is large, whereas most existing backdoor detection methods primarily focus on identifying smaller triggers.
@@ B.1 Baseline Attacks
1. KnockoffNet [4]: Following Tang et al. [11], we utilize KnockoffNet with a random strategy. The attacker randomly chooses samples from the surrogate dataset and gets the predictions of these samples by querying the target model. Subsequently, the attacker uses the sample-prediction pairs to train the surrogate model locally. In this paper, we set for all four dataset.
2. JBDA-TR [5]: JBDA-TR uses generated synthetic samples to test the decision boundaries of the target model. Specifically, JBDA-TR utilizes a small set of samples to query the target model and train a surrogate model locally (similar to KnockoffNet). In this paper, we set the initial size as for better performances. Subsequently, JBDA-TR performs adversarial attacks on the current query dataset and the current surrogate model following:
@@
1. S4L Attack [7]: The training loss function consists a CE loss and a semi-supervised loss, which helps train the model on both labeled and unlabeled data. The semi-supervised loss can be calculated as follows:
@@
2. Smoothing Attack [9]: Each sample is augmented and fed into the target model times, and the prediction is computed as the average of queries. In this paper, we set for all experiments with Smoothing Attack.
3. D-DAE [10]: Attackers train a defense detection model and a label recover model to detect and bypass potential defenses. In the default configuration, we use recover models trained on different output perturbation methods following [11]. For advanced attacks (oracle attack in Sec. 5.3), we train the recover model on different watermarking strategies. Specifically, D-DAE trains a 3-layer neural network to remove perturbations added by defenders, and the training dataset contains samples generated from small shadow models trained on public datasets.
4. p-Bayes Attack [11]: Attackers use independent and neighborhood sampling to perform Bayes-based estimation for original labels. The attacker builds a look-up table:
@@ B.2 Baseline Defenses
1. EWE [16]: Defenders utilize the Soft Nearest Neighbor Loss (SNNL) to minimize the distance between watermark features and natural features. The SNNL can be calculated as:
@@
2. MEA-Defender [20]: Defenders introduce the utility loss, the watermarking loss, and the evasion loss to balance the model availability and watermark success rate. The training loss function can be represented as:
@@ B.3 Parameters
Hyperparameters for training target and surrogate models. Following [11], we use models trained on ImageNet [28] to initialize the target model. Tab. XI and Tab. XII list the hyperparameters used while training target and surrogate models.
@@
Hyperparameters of Neural Honeytrace on different datasets. Tab. XIII lists the hyperparameters of Neural Honeytrace when watermarking different models trained on different target datasets.

### v4

@@ A.1 Performance Comparison
A.2 Overhead of Different Defenses
Fig. 9 compares the defense overhead of different triggerable watermarking methods for protecting target models trained on various datasets. During the training phase, previous triggerable watermarking methods introduce additional time costs due to modifications in the training process. For example, EWE(Jia et al., 2021) adds regularization terms to the loss function, increasing the time cost of each forward-backward pass. Composite Backdoor (Lin et al., 2020) and MEA-Defender (Lv et al., 2024b) expand the training dataset with generated data. In contrast, Neural Honeytrace is training-free and can be directly applied to the target model without training.
A.3 Additional Ablation Study
Different Query Datasets. In Tab. 9, we compare the performance of Neural Honeytrace on a target model trained on CIFAR-100 when attackers use CIFAR-10, CIFAR-100, and TinyImageNet as surrogate datasets, respectively. The experimental results show that for out-of-distribution surrogate datasets (CIFAR-10 and TinyImageNet), attackers achieve similar extraction accuracy, and Neural Honeytrace maintains high watermark success rates. However, when using the same training dataset as the surrogate (CIFAR-100), attackers achieve higher extraction accuracy, while the watermark success rate of decreases.
The other columns in Fig. 10 show the effectiveness of the three hyperparameters, , of Neural Honeytrace. These hyperparameters are used to balance the model availability and the watermark success rate. Intuitively, larger and smaller leads to stronger watermarks but lower protected accuracy, which can also be observed in Fig. 10. Therefore, given a test dataset and the acceptable maximum drop in accuracy, model owners can identify appropriate hyperparameters.
Test-time Backdoor Detection. Fig. 11 presents additional results evaluating existing watermarking strategies against the test-time backdoor detection method BTI(Tao et al., 2022). As depicted, BTI fails to effectively detect watermarks embedded in surrogate models for several reasons: (1) For EWE, the watermark success rate is significantly lower compared to conventional backdoor attacks, making it challenging to search for triggers. (2) For MEA-Defender, the composite trigger’s size is large, whereas most existing backdoor detection methods primarily focus on identifying smaller triggers.
| Method | Naive | Oracle | ||||
| Acc | Fidelity | WSR | Acc | Fidelity | WSR | |
| EWE | 88.71% | 88.18% | 39.90% | 86.02% | 87.95% | 4.40% |
| Composite | 85.47% | 87.41% | 43.80% | 85.25% | 87.63% | 10.40% |
| MEA-Defender | 88.15% | 88.63% | 61.80% | 86.10% | 87.95% | 10.80% |
| Neural H/T | 83.23% | 81.16% | 65.00% | 64.19% | 65.28% | 11.80% |
Oracle Attack. We also evaluate Neural Honeytrace against a highly capable attacker with full knowledge of the defense mechanism. Specifically, this attacker possesses a set of samples containing both the ground-truth predictions made by the target model and the corresponding modified predictions generated by watermarking strategies. Using these prediction pairs, the attacker trains a label recovery network following the method in (Chen et al., 2023) and subsequently trains a surrogate model using the recovered predictions.
As shown in Tab. 10, the watermark success rate of triggerable watermarking methods decreases significantly under oracle attacks compared to Naive Attack. However, the extraction accuracy and fidelity achieved by oracle attacks against Neural Honeytrace are substantially lower than those against previous methods. Furthermore, compared to perturbation-based defenses(Tang et al., 2024), Neural Honeytrace demonstrates even lower extraction accuracy and fidelity, indicating that the stolen model fails to accurately reconstruct the functionality of the target model. Therefore, Neural Honeytrace is more robust than existing methods when defending against oracle attacks. More importantly, model owners can easily change the watermarking details since Neural Honeytrace supports plug-and-play implementation, making it more difficult for attackers to access the defense information.
A.4 Neural Honeytrace for Generative Models
Compared to classification models, the output space of generative models such as GAN and diffusion models has higher dimensions. According to our watermark transmission model, larger output spaces lead to better watermark transmission. Therefore, we evaluate Neural Honeytrace for Denoising Diffusion Probabilistic Models (Ho et al., 2020) (DDPM) trained on CIFAR-10. DDPM trains a U-Net model to predict the noise added in the current step and generates an image by iterative denoising:
@@ B.1 Baseline Attacks
1. KnockoffNet (Orekondy et al., 2019): Following Tang et al. (Tang et al., 2024), we utilize KnockoffNet with a random strategy. The attacker randomly chooses samples from the surrogate dataset and gets the predictions of these samples by querying the target model. Subsequently, the attacker uses the sample-prediction pairs to train the surrogate model locally. In this paper, we set for all four datasets.
2. JBDA-TR (Juuti et al., 2019): JBDA-TR uses generated synthetic samples to test the decision boundaries of the target model. Specifically, JBDA-TR utilizes a small set of samples to query the target model and train a surrogate model locally (similar to KnockoffNet). In this paper, we set the initial size as for better performances. Subsequently, JBDA-TR performs adversarial attacks on the current query dataset and the current surrogate model following:
@@
1. S4L Attack (Jagielski et al., 2020): The training loss function consists a CE loss and a semi-supervised loss, which helps train the model on both labeled and unlabeled data. The semi-supervised loss can be calculated as follows:
@@
2. Smoothing Attack (Lukas et al., 2022): Each sample is augmented and fed into the target model times, and the prediction is computed as the average of queries. In this paper, we set for all experiments with Smoothing Attack.
3. D-DAE (Chen et al., 2023): Attackers train a defense detection model and a label recover model to detect and bypass potential defenses. In the default configuration, we use recover models trained on different output perturbation methods following (Tang et al., 2024). For advanced attacks, we train the recover model on different watermarking strategies. Specifically, D-DAE trains a 3-layer neural network to remove perturbations added by defenders, and the training dataset contains samples generated from small shadow models trained on public datasets.
4. p-Bayes Attack (Tang et al., 2024): Attackers use independent and neighborhood sampling to perform Bayes-based estimation for original labels. The attacker builds a look-up table:
@@ B.2 Baseline Defenses
1. EWE (Jia et al., 2021): Defenders utilize the Soft Nearest Neighbor Loss (SNNL) to minimize the distance between watermark features and natural features. The SNNL can be calculated as:
@@
2. MEA-Defender (Lv et al., 2024b): Defenders introduce the utility loss, the watermarking loss, and the evasion loss to balance the model availability and watermark success rate. The training loss function can be represented as:
@@ B.3 Parameters
Hyperparameters for training target and surrogate models. Following (Tang et al., 2024), we use models trained on ImageNet (Deng et al., 2009) to initialize the target model. Tab. 11 and Tab. 12 list the hyperparameters used while training target and surrogate models.
@@
Hyperparameters of Neural Honeytrace on different datasets. Tab. 13 lists the hyperparameters of Neural Honeytrace when watermarking different models trained on different target datasets.